The value of time for 20 oscillations are obtained 16.1 15.9 16.0 16.2 15.8 Determine the period of the pendulum


1
Expert's answer
2021-04-01T07:50:27-0400

Let's first find the average time for 20 oscillations:


"t_{avg}=\dfrac{16.1\ s+15.9\ s+16.0\ s+16.2\ s+15.8\ s}{5}=16\ s."

Finally, we can find the period of the pendulum:


"T=\dfrac{t_{avg}}{20}=\dfrac{16\ s}{20}=0.8\ s."